OpenStudy (anonymous):

suppose y varies directly with x and y=25 when x=140 what is the value of x when y=36

OpenStudy (anonymous):

x = 126

OpenStudy (anonymous):

x+y = x+y 140+25 = x + 36 165 = x+36 165-36 = x 126 = x

OpenStudy (shubhamsrg):

that is the wrong solution @lindie y varies directly with x means y is directly proportional to x, or y = kx , for some constant k.